You deposit $500 into a savings account. You know that the future value will be $800 in four years
time. What is the annual interest rate you receive from the bank?
Answer #1

We are given

Present value of deposit=PV=$500

Future value of deposit=FV=$800

Time duration=n=4 years

Let the annual interest rate be i

We know that future value of a single cash flow is given by

FV=PV*(1+i)n

800=500*(1+i)4

(1+i)=(800/500)1/4=1.124683

i=0.124683 or 12.47%